# Efficient Neuro-Symbolic Retrieval-Augmented Generation through Adaptive Query Routing

### v1

@@ 0.A.7 Ablation Studies & Adaptive Logic Criticality
0.A.8 Qualitative Analysis Examples
This section demonstrates SymRAG’s reasoning capabilities through trace examples, illustrating the interplay between symbolic and neural components using color-coded analysis where blue represents correct answers/outputs, green indicates evidence and correct processing steps, and red denotes potential errors or alternative paths.
Table 5: Combined Complex Multi-Hop and Discrete Reasoning Traces ("CA" = Correct Answer)
@@ 0.A.8 Qualitative Analysis Examples
Path Selection: Given $\kappa>T_{\mathrm{neural}}=0.7$ and moderate resource pressure, Algorithm 1 selects hybrid path
@@
Information Extraction: Neural processing encounters conflicting birth years: 1951 (secondary source) vs 1952 (verified sources), identifies "Forrest Gump" as Best Picture winner and "Robert Zemeckis" as director
Hybrid Integration: Alignment Layer (see Appendix 0.A.5) resolves conflicts through cross-reference validation, rejecting inconsistent birth year, producing confidence-weighted answer: 1952
@@
Answer Synthesis: Structured fusion (see Appendix 0.A.5) combines counts: 3 touchdowns – 1 field goal = 2
@@
Fusion Validation: Both symbolic and neural components identified the same entities and temporal scope, with exact numerical agreement triggering confidence boost according to Definition 5, demonstrating the effectiveness of the reconciliation strategy.
0.A.8.1 Component Interaction Analysis
@@
Confidence-Driven Integration: The fusion confidence mechanism (Definition 5) successfully identifies high-confidence scenarios ($C_{fusion}=0.94$, $0.92$) where both components agree, enabling reliable answer selection without manual intervention.
0.A.9 Baseline Comparison & Performance Analysis
@@
Model System EM (%) F1 (%) Avg Time (s) CPU (%) GPU (%) Mem (%) Llama-3.2-3B SymRAG (Full) 99.4 89.4 0.985 4.6 41.1 5.6 Neural-Only 97.8 87.8 0.904 4.8 44.3 5.8 Symbolic-Only 31.4* 31.8 0.362 5.1 25.3 6.2 Mistral-7B SymRAG (Full) 97.6 91.8 2.443 3.6 66.0 8.0 Neural-Only 95.2 89.4 2.241 4.2 72.8 8.3 Symbolic-Only 31.4 31.8 0.362 5.1 25.3 6.2 * $p<0.05$ Statistical significance via paired t-tests with Bonferroni correction . Best results in bold.
@@
0.A.10 Statistical Analysis & Implementation Details
@@
0.A.10.1 Hardware & Software Configuration
@@
0.A.11 Related Work Comparison
@@
SymRAG HotpotQA: 100% EM (+4.7pp)a; DROP: 99.4% EM (+11.6pp)b 4.6% CPU / 41.1% GPUc Dynamic routing (PS/PN/PH)d Optimal pipeline per query & load
CDF-RAG [5] MedMCQA: 94.2% EM (+16.0pp)e Not quantified RL-based query refinement Static retrieval; high RL-loop latency Controller adds $\sim$7ms vs. $\sim$100ms RL loopf
CRP-RAG [9] HotpotQA: F1 +7.4ppg Not quantified Graph-guided path selection Multiple sequential LLM calls 100% EM with single LLM passh
RuleRAG [1] RuleQA: EM +103%i Not quantified Rule-ICL & fine-tuning Static rules; brittle under load Invokes KG only when gain $>$10%j
RuAG [11] DWIE: 92.6% F1 (+8.3pp)k Not quantified MCTS-based rule distillation Prompt-only rules; no retrieval boost Hybrid fusion of KG & neural context
HybridRAG [8] Nifty-50 Q&A: Recall@k +30%l Not quantified None (static fusion) Cannot adapt when one modality degrades Dynamically routes around underperforming modalities
@@
    PS/PN/PH: Symbolic/Neural/Hybrid paths (Alg. 1).
@@
    $\Delta$ vs. prior SOTA (78.2% EM) on MedMCQA [5].
@@
    $\Delta$ vs. baseline F1 (67.3%) on HotpotQA [9].
@@
    Relative EM improvement over 5 RuleQA benchmarks [1].
@@
    $\Delta$ vs. baseline F1 (84.3%) on DWIE [11].
@@
    $\Delta$ vs. vector-only Recall@k (+23pp) on Nifty-50 [8]. See Sec. 4 for full discussion of baselines and measurements.
0.A.12 System Complexity & Performance Optimization
@@ 0.A.13 Future Directions
0.A.14 Limitations & Ethical Considerations

### v2

@@ A.7 Ablation Studies & Adaptive Logic Criticality
A.8 Qualitative Analysis Examples
This section demonstrates SymRAG’s reasoning capabilities through trace examples, illustrating the interplay between symbolic and neural components using color‐coded analysis where blue represents correct answers/outputs, green indicates evidence and correct processing steps, and red denotes potential errors or alternative paths (Table 5).
Table 5: Combined Complex Multi-Hop and Discrete Reasoning Traces (“CA” = Correct Answer)
@@ A.8 Qualitative Analysis Examples
Path Selection: Given $\kappa>T_{\mathrm{neural}}=0.7$ and moderate resource pressure, Algorithm LABEL:alg:path-selection selects hybrid path
@@
Information Extraction: Neural processing encounters conflicting birth years: 1951 (secondary source) vs 1952 (verified sources), identifies “Forrest Gump” as Best Picture winner and “Robert Zemeckis” as director
Hybrid Integration: Alignment Layer (see Appendix A.5) resolves conflicts through cross-reference validation, rejecting inconsistent birth year, producing confidence-weighted answer: 1952
@@
Answer Synthesis: Structured fusion (see Appendix A.5) combines counts: 3 touchdowns – 1 field goal = 2
@@
Fusion Validation: Both symbolic and neural components identified the same entities and temporal scope, with exact numerical agreement triggering confidence boost according to Definition A.3, demonstrating the effectiveness of the reconciliation strategy.
A.8.1 Component Interaction Analysis
@@
Confidence‐Driven Integration: The fusion confidence mechanism (Definition A.3) successfully identifies high‐confidence scenarios ($C_{fusion}=0.94$, $0.92$) where both components agree, enabling reliable answer selection without manual intervention.
A.9 Baseline Comparison & Performance Analysis
@@
Model System EM (%) F1 (%) Avg Time (s) CPU (%) GPU (%) Mem (%) Llama-3.2-3B SymRAG (Full) 99.4 89.4 0.985 4.6 41.1 5.6 Neural-Only 97.8 87.8 0.904 4.8 44.3 5.8 Symbolic-Only 31.4* 31.8 0.362 5.1 25.3 6.2 Mistral-7B SymRAG (Full) 97.6 91.8 2.443 3.6 66.0 8.0 Neural-Only 95.2 89.4 2.241 4.2 72.8 8.3 Symbolic-Only 31.4 31.8 0.362 5.1 25.3 6.2 * $p<0.05$ Statistical significance via paired t‐tests with Bonferroni correction . Best results in bold.
@@
A.10 Statistical Analysis & Implementation Details
@@
A.10.1 Hardware & Software Configuration
@@
A.11 Related Work Comparison
@@
SymRAG HotpotQA: 100% EM a; DROP: 99.4% EM b 4.6% CPU 41.1% GPUc Dynamic routing (PS/PN/PH)d Optimal pipeline per query & load
CDF-RAG (Khatibi et al., 2025) MedMCQA: 94.2% EM (+16.0pp)e Not quantified RL-based query refinement Static retrieval; high RL-loop latency Controller adds $\sim$7ms vs. $\sim$100ms RL loopf
CRP-RAG (Xu et al., 2024) HotpotQA: F1 +7.4ppg Not quantified Graph-guided path selection Multiple sequential LLM calls 100% EM with single LLM passh
RuleRAG (Chen et al., 2024) RuleQA: EM +103%i Not quantified Rule-ICL & fine-tuning Static rules; brittle under load Invokes KG only when gain $>$10%j
RuAG (Zhang et al., 2024) DWIE: 92.6% F1 (+8.3pp)k Not quantified MCTS-based rule distillation Prompt-only rules; no retrieval boost Hybrid fusion of KG & neural context
HybridRAG (Sarmah et al., 2024) Nifty-50 Q&A: Recall@k +30%l Not quantified None (static fusion) Cannot adapt when one modality degrades Dynamically routes around underperforming modalities
@@
    PS/PN/PH: Symbolic/Neural/Hybrid paths (Alg. LABEL:alg:path-selection).
@@
    $\Delta$ vs. prior SOTA (78.2% EM) on MedMCQA (Khatibi et al., 2025).
@@
    $\Delta$ vs. baseline F1 (67.3%) on HotpotQA (Xu et al., 2024).
@@
    Relative EM improvement over 5 RuleQA benchmarks (Chen et al., 2024).
@@
    $\Delta$ vs. baseline F1 (84.3%) on DWIE (Zhang et al., 2024).
@@
    $\Delta$ vs. vector-only Recall@k (+23pp) on Nifty-50 (Sarmah et al., 2024). See Sec. 4 for full discussion of baselines and measurements.
A.12 System Complexity & Performance Optimization
@@ A.13 Future Directions
A.14 Limitations & Ethical Considerations